# DogFit: Domain-guided Fine-tuning for Efficient Transfer Learning of Diffusion Models

### v6

@@ Abstract
Codehttps://github.com/yaramohamadi/DogFit
@@ Introduction
Denoising diffusion probabilistic models (DDPMs) (sohl2015deep; ho2020denoising) have emerged as powerful generative models, achieving state-of-the-art (SOTA) results in image synthesis (dhariwal2021diffusion; rombach2022high), video generation (gupta2024photorealistic; ho2022video), and image editing (kawar2023imagic). However, producing high-quality and diverse outputs still requires substantial computational and data resources, particularly with large diffusion backbones and image sizes. In practical applications such as personalized generation with limited data, full model training from scratch is infeasible, making transfer learning a critical tool for target adaptation. A common approach in this setting is to fine-tune a pre-trained diffusion model on the target data. While straightforward, this strategy frequently suffers from overfitting and poor generalization due to the limited size and diversity of the target domain.
@@
Figure 1: DogFit outperforms SOTA guidance methods in majority of cases for diffusion transfer learning, without increasing sampling overhead (TFLOPS). Evaluation is based on FID and $\text{FD}_{\text{DINOV2}}$ averaged across six target datasets.
@@
Classifier-free guidance (CFG) is a standard technique for improving generation quality at inference time by modulating the conditional signal (ho2022classifier; karras2024guiding). This technique is often used in its original form even after target adaptation to steer the model towards better generalization. However, CFG assumes access to well-trained conditional and unconditional models, which is difficult to ensure in transfer scenarios where the target domain is small or lacks labels (ho2022classifier; zhong2025domain). Recent methods like domain guidance (DoG) address these issues by utilizing the unconditional source model as a stronger marginal noise estimator, offering improved target domain alignment and generalization  (zhong2025domain; phunyaphibarn2025unconditional). Nevertheless, both CFG and DoG introduce significant computational overhead at sampling time due to their reliance on dual forward passes (ho2022classifier; zhong2025domain). On the other hand, model guidance (MG) mitigates CFG’s runtime cost in in-domain settings by injecting the guidance signal directly into the diffusion training objective (chen2025visual; tang2025diffusion). However, MG inherits the limitations of CFG in transfer learning due to the underfitting of the unconditional noise estimator (zhong2025domain). Further, MG hard-codes the guidance strength value during training, restricting control at inference. These limitations raise a key question: can we design effective guidance mechanisms for diffusion transfer learning without incurring computational overhead and still maintaining controllability over the guidance strength?
@@
This paper proposes domain-guided fine-tuning (DogFit), a method that injects a domain-aware guidance offset into the training loss during the fine-tuning process. This allows the diffusion model to directly learn the guided direction, removing the need for additional test-time processes. The domain-aware design is motivated by our observation that during fine-tuning, the source model offers stronger marginal noise estimates than the evolving target model. This design encourages the model to step toward the target domain manifold and avoid out-of-distribution generation. To support efficient control over the guidance strength at inference, DogFit conditions the model on the guidance value and exposes it to a range of such values during training. This allows the model to modulate its generation behavior accordingly, enabling fidelity–diversity trade-offs at inference with negligible sampling overhead. We further investigate the optimal placement and timing of the guidance offset during training. Two cost-effective guidance scheduling mechanisms are proposed for DogFit: (1) a late-start strategy that delays guidance until the model has learned sufficiently stable target representations; and (2) a cut-off scheme that restricts guidance to the later denoising steps, where fine-grained domain-specific details are more prevalent. Results over six datasets with different distribution shifts and supervision levels, using two SOTA diffusion backbones, DiT (peebles2023scalable) and SiT (ma2024sit), show that DogFit can outperform SOTA guidance methods in terms of FID and $\text{FD}_{\text{DINOV2}}$ while reducing their sampling TFLOPS by up to 2× (Fig.1). Results establish DogFit as a practical and efficient guidance method for diffusion transfer learning.
@@
Efficient Diffusion Models. The iterative denoising process in diffusion models imposes high sampling-time costs (shen2025efficient). To accelerate generation, prior work reduces the number of sampling steps via improved solvers (song2020denoising; lu2022dpm), optimized noise schedules (zheng2023improved), or distillation-based methods that train few(one)-step student models to mimic large multi-step teachers (salimans2022progressive; yin2024one). Recent efforts  (jensen2025efficient; hsiao2024plug; zhou2025dice) distill CFG to eliminate test-time dual passes. Our method, similar to CFG distillation, aims to internalize guidance, but does so without the architectural modifications or post-hoc training stages required by distillation methods. Instead, DogFit simply integrates guidance directly into the fine-tuning objective.
@@
Diffusion Transfer Learning. Transfer learning adapts a model pre-trained on a large-scale source domain to a target domain with smaller size and diversity, leveraging learned representations for better generalization (weiss2016survey). In the context of diffusion models, strategies broadly fall into three categories: Parameter-efficient fine-tuning (PEFT) methods reduce training cost by limiting the number of trainable parameters, often via adapters or LoRA (hu2022lora; xie2023difffit; moon2022fine). Distillation-based methods preserve pre-trained source priors during adaptation, particularly during early denoising steps (zhong2024diffusion; hur2024expanding). Few-shot image generation methods enable diffusion models to generalize to unseen domains using a handful of images (zhu2022few; wang2024bridging; cao2024few; ouyang2024transfer; ruiz2023dreambooth). Despite this progress, current methods overlook the role of guidance in shaping the generation trajectory, implicitly assuming that its original form remains optimal after adaptation, a limiting assumption in small, low-diversity target domains. We propose a new guidance method tailored for diffusion transfer that can be layered on top of existing fine-tuning methods for improving generalization without incurring additional sampling cost.
@@
Diffusion Models. Denoising Diffusion Probabilistic Models (DDPMs) (ho2020denoising) are generative models that learn to reverse a fixed noising process applied over $T$ time-steps. Starting from a clean image $x_{0}$, noise is gradually added to produce a sequence of noisy images $\{x_{t}\}_{t=1}^{T}$. A neural network $\epsilon_{\theta}(x_{t},t)$ is trained to predict the noise $\epsilon$ at each time-step $t$, using this denoising objective:
@@
Time-step $t$ is omitted in subsequent sections when clear from context. In conditional generation, the model receives an auxiliary input $c$ (e.g., class labels or text), producing $\epsilon_{\theta}(x_{t}|c)$. SOTA transformer-based diffusion architectures such as DiT (peebles2023scalable) and SiT (ma2024sit) operate in latent space rather than directly on pixel space. For notational simplicity, however, we keep using $x$ to denote the model input.
Guidance in Transfer Learning. While diffusion models can functionally incorporate conditional inputs during training, their outputs often drift under weak conditioning, motivating the use of guidance signals during the reverse process  (dhariwal2021diffusion; ho2022classifier). We present three representative guidance strategies in transfer learning (i.e., CFG, MG, and DoG) covering the main design axes: where the guidance signal originates (unconditional branch, or external source model) and when it is injected (sampling-time vs. training-time). See visual comparison in Fig. 2.
@@
- Classifier-Free Guidance (CFG) (ho2022classifier; nichol2021glide) improves conditional generation by extrapolating the model’s class-conditional prediction further away from its unconditional counterpart-typically using the same model. The reverse process is modified as:
@@
where increasing the guidance scale $w>1$ amplifies the effect of the condition during sampling. CFG poses limitations in transfer settings: The unconditional noise estimator is prone to underfitting due to joint training under limited target training data (chen2023score). Further, CFG often steers generations toward out-of-distribution regions in the target domain (Fig. 2 (b)), and it relies on conditionally labeled data, which may be unavailable in some domains (zhong2025domain).
- Domain Guidance (DoG) (zhong2025domain) offers a stronger alternative in transfer learning by using the unconditional source model as the marginal noise estimator:
@@
where $\boldsymbol{\epsilon}_{\theta}(x_{t}\mid c,\mathcal{D^{T}})$ is the fine-tuned target model’s class-conditional prediction and $\boldsymbol{\epsilon}_{\theta_{0}}(x_{t})$ is the unconditional one of the source model. The source model, having been trained on rich data, provides a stronger reference point than CFG and further improves target domain alignment (zhong2025domain) (Fig. 2 (c)). Like CFG, however, DoG requires two forward passes during sampling.
- Model Guidance (MG) (tang2025diffusion) removes the need for CFG-style computation during sampling by incorporating its effect directly into the training objective. The model is trained to match a guidance-enhanced noise target:
@@
+ CFG (ho2022classifier) 20.83 24.07 5.03 11.77 10.60 14.46 x2 732.28
+ DoG (zhong2025domain) 13.87 17.30 23.76 3.65 10.97 9.77 13.09 x2 732.28
MG (tang2025diffusion) 19.91 23.71 4.85 11.13 11.03 14.13 x1 366.14
@@
+ CFG (ho2022classifier) 314.32 401.96 200.92 410.48 227.46 311.03 x2 732.28
+ DoG (zhong2025domain) 273.93 266.25 377.49 135.74 314.16 132.91 245.31 x2 732.28
MG (tang2025diffusion) 299.06 409.20 220.40 379.39 255.88 312.78 x1 366.14
@@
+ CFG (ho2022classifier) - 9.28 23.21 3.49 7.95 9.71 10.73 x2 1083.77
+ DoG (zhong2025domain) 7.63 11.22 23.53 3.39 7.95 9.71 11.16 x2 1083.77
MG (tang2025diffusion) - 9.64 23.10 3.60 6.84 8.62 10.36 x1 454.01
@@
+ CFG (ho2022classifier) - 236.13 406.63 203.47 368.26 190.24 280.94 x2 1083.77
+ DoG (zhong2025domain) 335.15 215.12 387.89 160.85 298.65 139.09 240.32 x2 1083.77
MG (tang2025diffusion) - 232.17 418.86 229.32 359.90 203.23 288.70 x1 454.01
@@
[…]
[Uncaptioned image] Domain Guided Fine-Tuning
@@
The training architecture is illustrated in Fig. 3(c). The training algorithm is shown in Appx. The rest of this subsection provides theoretical insights motivating the design of DogFit (full statements and proofs in Appx.)
@@
where $\epsilon_{\theta}(x_{t}\mid c,\boldsymbol{1},\mathcal{D^{T}})$ is the unguided class-conditional prediction. During training, the model must learn to operate across a range of guidance strengths to support controllable fidelity–diversity trade-offs at inference, while still preserving stable behavior in the unguided setting. Our initial experiments revealed that a simple uniform sampling of $\boldsymbol{w}$ during training can reduce diversity and destabilize generation, even at lower guidance levels due to excessive exposure to high $w$. We sample $\boldsymbol{w}$ from a shifted exponential decaying distribution (SEDD; see distribution CDF in Appx.):
@@
where $\lambda$ controls the decay rate. This distribution favors smaller guidance strengths while occasionally introducing larger values, enabling guidance robustness without destabilizing training. As a result, the model generalizes well across the guidance spectrum and supports dynamic controllability at inference (Fig. 5; see comparison with uniform scheduling in Appx.). The next key question is how to actually incorporate $w$ into the model?
@@
Inspired by test-time guidance scheduling for improved generation quality (sadat2023cads; kynkaanniemi2024applying), we investigate the optimal timing and placement of training-time guidance using two simple yet effective strategies. First, we introduce a late-start mechanism that delays the injection of guidance until iteration $s>\tau_{\text{s}}$, avoiding noisy updates and unstable gradients in the early phases of training. This is particularly beneficial in conditional settings where class embeddings are reinitialized to accommodate mismatched source and target label spaces. By deferring guidance, the model benefits from more stable target representations and condition embeddings, resulting in higher-fidelity updates. Second, we introduce a cut-off strategy that disables guidance at higher noise levels ($t>\tau_{\text{c}}$), concentrating domain supervision on the later denoising stages where adaptation to target-specific details is most effective (choi2022perception). This design preserves source-driven diversity in the early denoising steps and prevents the model from over-relying on the target domain to shape the global structure of the image. This becomes crucial when source and target domains share similar structural priors (e.g. FFHQ faces).
@@
Setup. DogFit is validated on six challenging datasets with diverse domain characteristics and label availability for a comprehensive assessment: Food101 (bossard2014food), Caltech101 (griffin2007caltech), CUB-200-201 (wah2011caltech), ArtBench10 (liao2022artbench), Stanford-Cars (krause20133d), and FFHQ256 (karras2019style). These datasets range from fine-grained species (wah2011caltech) to abstract art-styles (liao2022artbench) and face generation (karras2019style), covering a variety of distribution shifts in appearance, texture, and structure. We compare DogFit to standard fine-tuning, CFG, DoG, and MG. All methods are applied on DiT-XL/2 (peebles2023scalable) and SiT-XL/2 (ma2024sit) backbones pre-trained on ImageNet at $\times256$ resolution for 7M steps with Fréchet Inception Distance (FID) (heusel2017gans) scores of 2.27 (DiT) and 2.06 (SiT). We generate 10,000 images using 50 sampling steps (peebles2023scalable; xie2023difffit), fixing guidance strength to 1.5 for fair comparison, and fine-tune for 24,000 training steps  (zhong2025domain) with batch size of 32 on 2 A100 GPUs for 5 hours. We calculate FID and $\text{FD}_{\text{DINOV2}}$ (stein2023exposing) between the generated images and the full datasets to evaluate the generation quality. Further, we compute Precision and Recall (kynkaanniemi2019improved) for analyzing fidelity and diversity.
@@
Tab.1 summarizes our main quantitative findings. DogFit consistently performs comparatively or superior to prior methods in both FID and $\text{FD}_{\text{DINOV2}}$, while maintaining the efficiency of one-pass sampling. Unlike CFG and MG, DogFit remains effective in unconditional settings like FFHQ due to its class-agnostic domain-alignment direction. On SiT, CFG and MG slightly outperform DogFit in FID, but DogFit outperforms both CFG and MG on $\text{FD}_{\text{DINOV2}}$, underscoring the value of utilizing diverse metrics for comprehensive evaluation. Importantly, equipping DogFit with controllable guidance introduces no performance degradation while increasing the parameter count by only 1.33 million parameters (less than 2% of the initial network parameters), leaving no notable impact on sampling TFLOPS. Additional results on Precision and Recall are provided in Appx. Fig. 4 presents a visual comparison of generated samples across target domains on DiT. Guidance-based methods consistently outperform naïve fine-tuning by producing more realistic and coherent images. DogFit generates sharp, domain-relevant details on par with or better looking than DoG, while avoiding off-distribution artifacts occasionally seen with CFG and MG (e.g., distorted faces or tires). SiT qualitative results are provided in Appx.
@@
Controllable Guidance We study the effect of $\lambda$ for sampling the guidance scalar $w$ from the SED during training (Eq. 8). As shown in Fig. 5 (left), sparse exposure to higher guidance is sufficient for generalization and achieving a diversity-fidelity tradeoff. In contrast, excessive exposure to large guidance can harm the model’s ability to retain pretrained features, leading to instability and collapsed generations. Based on these findings, we adopt $\lambda=3$ (95% of sampled $w$s in $[1,2]$) for all experiments. Fig. 5 (right) illustrates how varying guidance strength at test time effectively modulates the trade-off between fidelity and diversity, increasing sample quality while reducing variation with higher guidance values (more analysis in Appx.
Guidance Scheduling. We analyze the impact of late-start and cut-off on generation quality and diversity. Fig. 6(a) shows how varying the late-start threshold $\tau_{s}$ impacts FID, Precision, and Recall. Delaying guidance improves FID and especially Precision at first, as early stabilization allows formation of more accurate directional cues. However, setting $\tau_{s}$ too high leaves insufficient time for the model to internalize guidance, reducing it to naïve fine-tuning and lowering precision. We set $\tau_{s}=12\text{k}$ for FID and $\tau_{s}=6\text{k}$ for $\text{FD}_{\text{DINOV2}}$ in all experiments. Fig. 6(b) shows the effect of cut-off threshold $\tau_{c}$ on FID, Precision, and Recall. Delaying guidance to later denoising steps enhances diversity, particularly for target datasets with strong global structural similarity to the source domain, such as FFHQ, Cars, and Caltech. However, setting $\tau_{c}$ too low limits exposure to guidance, again reducing the model to naïve fine-tuning. While the optimal $\tau_{c}$ may depend on the source–target similarity, we use $\tau_{c}=0.5$ for FID and $\tau_{c}=1$ for $\text{FD}_{\text{DINOV2}}$ in our experiments ($t\in[0,1]$). See $\text{FD}_{\text{DINOV2}}$ analysis in Appx.
# Steps Food FID↓ Art FID↓
@@
Table 2: Varying the number of sampling steps for MG and DogFit.
@@
Table 3: Using DogFit with PEFT.
Domain-guided Parameter-efficient Fine-tuning. We demonstrate that DogFit is fully compatible with DiffFit, a leading PEFT method for DiT models (xie2023difffit). As shown in Tab.3, combining the two yields high-quality generations with lower cost in both training and test-time. This highlights DogFit’s practicality as a simple drop-in enhancement for efficient fine-tuning mechanisms.

### v7

# Yara Bahram, Mohammadhadi Shateri, Eric Granger

@@ Abstract
Code is available at https://github.com/yaramohamadi/DogFit.
@@ Introduction
Denoising diffusion probabilistic models (DDPMs) (Sohl-Dickstein et al. 2015; Ho et al. 2020) have emerged as powerful generative models, achieving state-of-the-art (SOTA) results in image synthesis (Dhariwal and Nichol 2021; Rombach et al. 2022), video generation (Gupta et al. 2024; Ho et al. 2022), and image editing (Kawar et al. 2023). However, producing high-quality and diverse outputs still requires substantial computational and data resources, particularly with large diffusion backbones and image sizes. In practical applications such as personalized generation with limited data, full model training from scratch is infeasible, making transfer learning a critical tool for target adaptation. A common approach in this setting is to fine-tune a pre-trained diffusion model on the target data. While straightforward, this strategy frequently suffers from overfitting and poor generalization due to the limited size and diversity of the target domain.
@@
Figure 1: DogFit achieves competitive average FID and $\mathrm{FD}_{\text{DINOV2}}$ when compared to SOTA guidance mechanisms for diffusion transfer learning, all without increasing the computational complexity (sampling TFLOPS). Averages are computed over six target datasets.
@@
Classifier-free guidance (CFG) is a standard technique for improving generation quality at inference time by modulating the conditional signal (Ho and Salimans 2022; Karras et al. 2024). This technique is often used in its original form even after target adaptation to steer the model towards better generalization. However, CFG assumes access to well-trained conditional and unconditional models, which is difficult to ensure in transfer scenarios where the target domain is small or lacks labels (Ho and Salimans 2022; Zhong et al. 2025). Recent methods like domain guidance (DoG) address these issues by utilizing the unconditional source model as a stronger marginal noise estimator, offering improved target domain alignment and generalization  (Zhong et al. 2025; Phunyaphibarn et al. 2025). Nevertheless, both CFG and DoG introduce significant computational overhead at sampling time due to their reliance on dual forward passes (Ho and Salimans 2022; Zhong et al. 2025). On the other hand, model guidance (MG) mitigates CFG’s runtime cost in in-domain settings by injecting the guidance signal directly into the diffusion training objective (Chen et al. 2025; Tang et al. 2025). However, MG inherits the limitations of CFG in transfer learning due to the underfitting of the unconditional noise estimator (Zhong et al. 2025). Further, MG hard-codes the guidance strength value during training, restricting control at inference. These limitations raise a key question: can we design effective guidance mechanisms for diffusion transfer learning without incurring computational overhead and still maintaining controllability over the guidance strength?
@@
This paper proposes domain-guided fine-tuning (DogFit), a method that injects a domain-aware guidance offset into the training loss during the fine-tuning process. This allows the diffusion model to directly learn the guided direction, removing the need for additional test-time processes. The domain-aware design is motivated by our observation that during fine-tuning, the source model offers stronger marginal noise estimates than the evolving target model. This design encourages the model to step toward the target domain manifold and avoid out-of-distribution generation. To support efficient control over the guidance strength at inference, DogFit conditions the model on the guidance value and exposes it to a range of such values during training. This allows the model to modulate its generation behavior accordingly, enabling fidelity–diversity trade-offs at inference with negligible sampling overhead. We further investigate the optimal placement and timing of the guidance offset during training. Two cost-effective guidance scheduling mechanisms are proposed for DogFit: (1) a late-start strategy that delays guidance until the model has learned sufficiently stable target representations; and (2) a cut-off scheme that restricts guidance to the later denoising steps, where fine-grained domain-specific details are more prevalent. Results over six datasets with different distribution shifts and supervision levels, using two SOTA diffusion backbones, DiT (Peebles and Xie 2023) and SiT (Ma et al. 2024), show that DogFit can outperform SOTA guidance methods in terms of FID and $\text{FD}_{\text{DINOV2}}$ while reducing their sampling TFLOPS by up to 2× (Fig.1). Results establish DogFit as a practical and efficient guidance method for diffusion transfer learning.
@@
Efficient Diffusion Models. The iterative denoising process in diffusion models imposes high sampling-time costs (Shen et al. 2025). To accelerate generation, prior work reduces the number of sampling steps via improved solvers (Song et al. 2020; Lu et al. 2022), optimized noise schedules (Zheng et al. 2023), or distillation-based methods that train few(one)-step student models to mimic large multi-step teachers (Salimans and Ho 2022; Yin et al. 2024). Recent efforts  (Jensen and Sadat 2025; Hsiao et al. 2024; Zhou et al. 2025) distill CFG to eliminate test-time dual passes. Our method, similar to CFG distillation, aims to internalize guidance, but does so without the architectural modifications or post-hoc training stages required by distillation methods. Instead, DogFit simply integrates guidance directly into the fine-tuning objective.
@@
Diffusion Transfer Learning. Transfer learning adapts a model pre-trained on a large-scale source domain to a target domain with smaller size and diversity, leveraging learned representations for better generalization (Weiss et al. 2016). In the context of diffusion models, strategies broadly fall into three categories: Parameter-efficient fine-tuning (PEFT) methods reduce training cost by limiting the number of trainable parameters, often via adapters or LoRA (Hu et al. 2022; Xie et al. 2023; Moon et al. 2022). Distillation-based methods preserve pre-trained source priors during adaptation, particularly during early denoising steps (Zhong et al. 2024; Hur et al. 2024). Few-shot image generation methods enable diffusion models to generalize to unseen domains using a handful of images (Zhu et al. 2022; Wang et al. 2024; Cao and Gong 2024; Ouyang et al. 2024; Ruiz et al. 2023). Despite this progress, current methods overlook the role of guidance in shaping the generation trajectory, implicitly assuming that its original form remains optimal after adaptation, a limiting assumption in small, low-diversity target domains. We propose a new guidance method tailored for diffusion transfer that can be layered on top of existing fine-tuning methods for improving generalization without incurring additional sampling cost.
@@
Diffusion Models. Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al. 2020) are generative models that learn to reverse a fixed noising process applied over $T$ time-steps. Starting from a clean image $x_{0}$, noise is gradually added to produce a sequence of noisy images $\{x_{t}\}_{t=1}^{T}$. A neural network $\epsilon_{\theta}(x_{t},t)$ is trained to predict the noise $\epsilon$ at each time-step $t$, using this denoising objective:
@@
Time-step $t$ is omitted in subsequent sections when clear from context. In conditional generation, the model receives an auxiliary input $c$ (e.g., class labels or text), producing $\epsilon_{\theta}(x_{t}|c)$. SOTA transformer-based diffusion architectures such as DiT (Peebles and Xie 2023) and SiT (Ma et al. 2024) operate in latent space rather than directly on pixel space. For notational simplicity, however, we keep using $x$ to denote the model input.
Guidance in Transfer Learning. While diffusion models can functionally incorporate conditional inputs during training, their outputs often drift under weak conditioning, motivating the use of guidance signals during the reverse process  (Dhariwal and Nichol 2021; Ho and Salimans 2022). We present three representative guidance strategies in transfer learning (i.e., CFG, MG, and DoG) covering the main design axes: where the guidance signal originates (unconditional branch, or external source model) and when it is injected (sampling-time vs. training-time). See visual comparison in Fig. 2.
@@
- Classifier-Free Guidance (CFG) (Ho and Salimans 2022; Nichol et al. 2021) improves conditional generation by extrapolating the model’s class-conditional prediction further away from its unconditional counterpart-typically using the same model. The reverse process is modified as:
@@
where increasing the guidance scale $w>1$ amplifies the effect of the condition during sampling. CFG poses limitations in transfer settings: The unconditional noise estimator is prone to underfitting due to joint training under limited target training data (Chen et al. 2023). Further, CFG often steers generations toward out-of-distribution regions in the target domain (Fig. 2 (b)), and it relies on conditionally labeled data, which may be unavailable in some domains (Zhong et al. 2025).
- Domain Guidance (DoG) (Zhong et al. 2025) offers a stronger alternative in transfer learning by using the unconditional source model as the marginal noise estimator:
@@
where $\boldsymbol{\epsilon}_{\theta}(x_{t}\mid c,\mathcal{D^{T}})$ is the fine-tuned target model’s class-conditional prediction and $\boldsymbol{\epsilon}_{\theta_{0}}(x_{t})$ is the unconditional one of the source model. The source model, having been trained on rich data, provides a stronger reference point than CFG and further improves target domain alignment (Zhong et al. 2025) (Fig. 2 (c)). Like CFG, however, DoG requires two forward passes during sampling.
- Model Guidance (MG) (Tang et al. 2025) removes the need for CFG-style computation during sampling by incorporating its effect directly into the training objective. The model is trained to match a guidance-enhanced noise target:
@@
+ CFG (Ho and Salimans 2022) 20.83 24.07 5.03 11.77 10.60 14.46 x2 732.28
+ DoG (Zhong et al. 2025) 13.87 17.30 23.76 3.65 10.97 9.77 13.09 x2 732.28
MG (Tang et al. 2025) 19.91 23.71 4.85 11.13 11.03 14.13 x1 366.14
@@
+ CFG (Ho and Salimans 2022) 314.32 401.96 200.92 410.48 227.46 311.03 x2 732.28
+ DoG (Zhong et al. 2025) 273.93 266.25 377.49 135.74 314.16 132.91 245.31 x2 732.28
MG (Tang et al. 2025) 299.06 409.20 220.40 379.39 255.88 312.78 x1 366.14
@@
+ CFG (Ho and Salimans 2022) - 9.28 23.21 3.49 7.95 9.71 10.73 x2 1083.77
+ DoG (Zhong et al. 2025) 7.63 11.22 23.53 3.39 7.95 9.71 11.16 x2 1083.77
MG (Tang et al. 2025) - 9.64 23.10 3.60 6.84 8.62 10.36 x1 454.01
@@
+ CFG (Ho and Salimans 2022) - 236.13 406.63 203.47 368.26 190.24 280.94 x2 1083.77
+ DoG (Zhong et al. 2025) 335.15 215.12 387.89 160.85 298.65 139.09 240.32 x2 1083.77
MG (Tang et al. 2025) - 232.17 418.86 229.32 359.90 203.23 288.70 x1 454.01
@@
Domain Guided Fine-Tuning
@@
The training architecture is illustrated in Fig. 3(c). The training algorithm is shown in Appx. A. The rest of this subsection provides theoretical insights motivating the design of DogFit (full statements and proofs in Appx. A).
@@
where $\epsilon_{\theta}(x_{t}\mid c,\boldsymbol{1},\mathcal{D^{T}})$ is the unguided class-conditional prediction. During training, the model must learn to operate across a range of guidance strengths to support controllable fidelity–diversity trade-offs at inference, while still preserving stable behavior in the unguided setting. Our initial experiments revealed that a simple uniform sampling of $\boldsymbol{w}$ during training can reduce diversity and destabilize generation, even at lower guidance levels due to excessive exposure to high $w$. We sample $\boldsymbol{w}$ from a shifted exponential decaying distribution (SEDD; see distribution CDF in Appx. A):
@@
where $\lambda$ controls the decay rate. This distribution favors smaller guidance strengths while occasionally introducing larger values, enabling guidance robustness without destabilizing training. As a result, the model generalizes well across the guidance spectrum and supports dynamic controllability at inference (Fig. 5; see comparison with uniform scheduling in Appx. A)
The next key question is how to actually incorporate $w$ into the model?
@@
Inspired by test-time guidance scheduling for improved generation quality (Sadat et al. 2023; Kynkäänniemi et al. 2024), we investigate the optimal timing and placement of training-time guidance using two simple yet effective strategies. First, we introduce a late-start mechanism that delays the injection of guidance until iteration $s>\tau_{\text{s}}$, avoiding noisy updates and unstable gradients in the early phases of training. This is particularly beneficial in conditional settings where class embeddings are reinitialized to accommodate mismatched source and target label spaces. By deferring guidance, the model benefits from more stable target representations and condition embeddings, resulting in higher-fidelity updates. Second, we introduce a cut-off strategy that disables guidance at higher noise levels ($t>\tau_{\text{c}}$), concentrating domain supervision on the later denoising stages where adaptation to target-specific details is most effective (Choi et al. 2022). This design preserves source-driven diversity in the early denoising steps and prevents the model from over-relying on the target domain to shape the global structure of the image. This becomes crucial when source and target domains share similar structural priors (e.g. FFHQ faces).
@@
Setup. DogFit is validated on six challenging datasets with diverse domain characteristics and label availability for a comprehensive assessment: Food101 (Bossard et al. 2014), Caltech101 (Griffin et al. 2007), CUB-200-201 (Wah et al. 2011), ArtBench10 (Liao et al. 2022), Stanford-Cars (Krause et al. 2013), and FFHQ256 (Karras et al. 2019). These datasets range from fine-grained species (Wah et al. 2011) to abstract art-styles (Liao et al. 2022) and face generation (Karras et al. 2019), covering a variety of distribution shifts in appearance, texture, and structure. We compare DogFit to standard fine-tuning, CFG, DoG, and MG. All methods are applied on DiT-XL/2 (Peebles and Xie 2023) and SiT-XL/2 (Ma et al. 2024) backbones pre-trained on ImageNet at $\times256$ resolution for 7M steps with Fréchet Inception Distance (FID) (Heusel et al. 2017) scores of 2.27 (DiT) and 2.06 (SiT). We generate 10,000 images using 50 sampling steps (Peebles and Xie 2023; Xie et al. 2023), fixing guidance strength to 1.5 for fair comparison, and fine-tune for 24,000 training steps  (Zhong et al. 2025) with batch size of 32 on 2 A100 GPUs for 5 hours. We calculate FID and $\text{FD}_{\text{DINOV2}}$ (Stein et al. 2023) between the generated images and the full datasets to evaluate the generation quality. Further, we compute Precision and Recall (Kynkäänniemi et al. 2019) for analyzing fidelity and diversity.
@@
Tab.1 summarizes our main quantitative findings. DogFit consistently performs comparatively or superior to prior methods in both FID and $\text{FD}_{\text{DINOV2}}$, while maintaining the efficiency of one-pass sampling. Unlike CFG and MG, DogFit remains effective in unconditional settings like FFHQ due to its class-agnostic domain-alignment direction. On SiT, CFG and MG slightly outperform DogFit in FID, but DogFit outperforms both CFG and MG on $\text{FD}_{\text{DINOV2}}$, underscoring the value of utilizing diverse metrics for comprehensive evaluation. Importantly, equipping DogFit with controllable guidance introduces no performance degradation while increasing the parameter count by only 1.33 million parameters (less than 2% of the initial network parameters), leaving no notable impact on sampling TFLOPS. Additional results on Precision and Recall are provided in Appx. B. Fig. 4 presents a visual comparison of generated samples across target domains on DiT. Guidance-based methods consistently outperform naïve fine-tuning by producing more realistic and coherent images. DogFit generates sharp, domain-relevant details on par with or better looking than DoG, while avoiding off-distribution artifacts occasionally seen with CFG and MG (e.g., distorted faces or tires). SiT qualitative results are provided in Appx. B.
@@
Controllable Guidance We study the effect of $\lambda$ for sampling the guidance scalar $w$ from the SED during training (Eq. 8). As shown in Fig. 5 (left), sparse exposure to higher guidance is sufficient for generalization and achieving a diversity-fidelity tradeoff. In contrast, excessive exposure to large guidance can harm the model’s ability to retain pretrained features, leading to instability and collapsed generations. Based on these findings, we adopt $\lambda=3$ (95% of sampled $w$s in $[1,2]$) for all experiments. Fig. 5 (right) illustrates how varying guidance strength at test time effectively modulates the trade-off between fidelity and diversity, increasing sample quality while reducing variation with higher guidance values (more analysis in Appx. B).
Guidance Scheduling. We analyze the impact of late-start and cut-off on generation quality and diversity. Fig. 6(a) shows how varying the late-start threshold $\tau_{s}$ impacts FID, Precision, and Recall. Delaying guidance improves FID and especially Precision at first, as early stabilization allows formation of more accurate directional cues. However, setting $\tau_{s}$ too high leaves insufficient time for the model to internalize guidance, reducing it to naïve fine-tuning and lowering precision. We set $\tau_{s}=12\text{k}$ for FID and $\tau_{s}=6\text{k}$ for $\text{FD}_{\text{DINOV2}}$ in all experiments. Fig. 6(b) shows the effect of cut-off threshold $\tau_{c}$ on FID, Precision, and Recall. Delaying guidance to later denoising steps enhances diversity, particularly for target datasets with strong global structural similarity to the source domain, such as FFHQ, Cars, and Caltech. However, setting $\tau_{c}$ too low limits exposure to guidance, again reducing the model to naïve fine-tuning. While the optimal $\tau_{c}$ may depend on the source–target similarity, we use $\tau_{c}=0.5$ for FID and $\tau_{c}=1$ for $\text{FD}_{\text{DINOV2}}$ in our experiments ($t\in[0,1]$). See $\text{FD}_{\text{DINOV2}}$ analysis in Appx. B.
Steps Food FID↓ Art FID↓
@@
Table 2: Sampling steps for MG and DogFit methods using DiT-XL/2.
@@
Table 3: Combining DogFit with DiffFit.
Domain-guided Parameter-efficient Fine-tuning. We demonstrate that DogFit is fully compatible with DiffFit, a leading PEFT method for DiT models (Xie et al. 2023). As shown in Tab.3, combining the two yields high-quality generations with lower cost in both training and test-time. This highlights DogFit’s practicality as a simple drop-in enhancement for efficient fine-tuning mechanisms.